# The two-dimensional Rashba-Holstein model

### v1

@@ III Results and discussion
Figure 7: The (a) CDW and (b) SC correlation ratios as functions of $|U|$, in the limit of the attractive Hubbard model, for the case of a pure Rashba hopping. (c) The extrapolations to the thermodynamic limit of the $\sqrt{S_{CDW}/\text{L}^{2}}$, which leads to the order parameter presented in panel (d). The critical point is at $|U_{c}|=5.0\pm 0.5$).
@@
We proceed by projecting onto the ground state, assuming Lorentz invariance and defining $\beta\propto L^{z}$, where $z=1$ is the dynamic critical exponent at the QCP  Assaad and Herbut (2013); Otsuka et al. (2016); Parisen Toldin et al. (2015). As before, the critical point is determined by the crossing of $R_{\text{CDW}}$ for different lattice sizes, as shown in Fig. 6 (a). In this case, we have fixed $\beta=L$ and $\omega_{0}/\alpha=2$, while varying $\lambda$, with clear crossings around $\lambda/\alpha\approx 2.62$. Following a similar procedure as that used for determining the critical temperatures, we extrapolate the crossings for the critical $\lambda_{c}$ to the thermodynamic limit, as presented in Fig. 6 (b), thereby determining the QCP. We repeated this analysis for other values of $\omega_{0}$ [vide Fig. 6 (b)], which allowed us to construct the ground state phase diagram of the Rashba-Holstein model in the limit of the Weyl semimetal cones (i.e. a pure Rashba hopping), depicted in Fig. 6 (c). Notice that, even for high frequency (e.g., $\omega/\alpha=16$) the ground state is either a staggered CDW or a WSM one, with no enhancement of SC or striped CDW phases.
@@
Finally, the previous results naturally prompt the question of whether the coexistence between CDW and SC occurs only in the limit $\omega_{0}\to\infty$ (symmetry-imposed), or if it can also manifest at finite frequency values (not symmetry-imposed). To address this issue, we fix $\omega_{0}/\alpha=16$ and $\beta=L$, while varying $\lambda$. The results are displayed in Fig. 8, for the response of (a) the staggered CDW structure factor and (b) the homogeneous $s$-wave SC structure factor, $S_{sc}$. While the former shows significant enhancement for $\lambda\gtrsim 4.5$, consistent with a charge ordered ground state, the latter exhibits only weak size dependence as $\lambda$ increases, indicating the absence of SC long-range order. This should be compared to the case of the attractive Hubbard model, shown in Figs. 8 (c) and (d), for $S_{\rm CDW}$ and $S_{sc}$, respectively. In stark contrast, both the CDW and SC structure factors exhibit significant finite-size dependence, suggesting their coexistence in this regime. Therefore, if CDW and SC coexist for finite $\omega_{0}$, it likely occurs at very large frequency values or very small temperatures.
@@ IV Conclusions
However, the features of the model change drastically in the limit of a pure Rashba hopping, i.e., when $t/\alpha=0$. In this limit, four Weyl cones appear at the half-filling, and quantum phase transitions are expected to occur at strong interactions. Specifically, we demonstrated that the CDW phase emerges for a finite quantum critical coupling $\lambda_{c}$, for a given phonon frequency. By varying $\omega_{0}$, the critical $\lambda_{c}$ also changes, leading to the phase diagram shown in Fig. 6. At finite phonon frequencies, we expect this transition to belong to the Ising Gross-Neveu universality class with $N_{f}=8$ two component Dirac cones Herbut et al. (2009); Huffman and Chandrasekharan (2020). In the antiadiabatic limit where $1/\omega_{0}\to 0$ the interaction corresponds to the attractive Hubbard model, with $|U|=\lambda$. In this limit, and as described in Sec. II.2 we expect an emergent SU(2)η symmetry the ties SC and CDW orders together. In this limit the transitions for the WSM to the massive phase corresponds to a SU(2) Gross-Neveu transtion with $N_{f}=8$ two component Dirac cones. This corresponds to the same universality class as that of the Hubbard model on the honeycomb lattice Assaad and Herbut (2013); Otsuka et al. (2016). These findings provide a step forward in understanding the interplay between SC and CDW phases in systems with spin-orbit coupling, and we hope they could shed light on the nature of these compounds.
@@ ACKNOWLEDGMENTS
The authors are grateful to the Brazilian Agencies Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Ensino Superior (CAPES), and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro, FAPERJ. J.F. thanks FAPERJ, Grant No. SEI-260003/019642/2022. S.A.S.J. thanks CNPq, grant No. 201000/2024-5. N.C.C. acknowledges support from FAPERJ Grant No. E-26/200.258/2023 - SEI-260003/000623/2023, and CNPq Grant No. 313065/2021-7. FA thanks the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490).

### v2

@@ III Results and discussion
Figure 7: The (a) CDW and (b) SC correlation ratios as functions of $|U|$, in the limit of the attractive Hubbard model, for the case of a pure Rashba hopping. (c) The extrapolations to the thermodynamic limit of the $\sqrt{S_{CDW}/\text{L}^{2}}$, which leads to the order parameter presented in panel (d). The critical point is at $|U_{c}|=5.0\pm 0.5$.
@@
We proceed by projecting onto the ground state, assuming Lorentz invariance and defining $\beta\propto L^{z}$, where $z=1$ is the dynamic critical exponent at the QCP  Assaad and Herbut (2013); Otsuka et al. (2016); Parisen Toldin et al. (2015). As before, the critical point is determined by the crossing of $R_{\text{CDW}}$ for different lattice sizes, as shown in Fig. 6 (a). In this case, we have fixed $\beta=L$ and $\omega_{0}/\alpha=2$, while varying $\lambda$, with clear crossings around $\lambda/\alpha\approx 2.62$. Following a similar procedure as that used for determining the critical temperatures, we extrapolate the crossings for the critical $\lambda_{c}$ to the thermodynamic limit, as presented in Fig. 6 (b), thereby determining the QCP. We repeated this analysis for other values of $\omega_{0}$ [see, e.g., Fig. 6 (b)], which allowed us to construct the ground state phase diagram of the Rashba-Holstein model in the limit of the Weyl semimetal cones (i.e. a pure Rashba hopping), depicted in Fig. 6 (c). Notice that, even for high frequency (e.g., $\omega/\alpha=16$) the ground state is either a staggered CDW or a WSM one, with no enhancement of SC or striped CDW phases.
@@
Finally, the previous results naturally prompt the question of whether the coexistence between CDW and SC occurs only in the limit $\omega_{0}\to\infty$ (symmetry-imposed), or if it can also manifest at finite frequency values (not symmetry-imposed). To address this issue, we fix $\omega_{0}/\alpha=16$ and $\beta=L$, while varying $\lambda$. The results are displayed in Fig. 8, for the response of (a) the staggered CDW structure factor and (b) the homogeneous $s$-wave SC structure factor, $S_{\rm SC}$. While the former shows significant enhancement for $\lambda\gtrsim 4.5$, consistent with a charge ordered ground state, the latter exhibits only weak size dependence as $\lambda$ increases, indicating the absence of SC long-range order. This should be compared to the case of the attractive Hubbard model, shown in Figs. 8 (c) and (d), for $S_{\rm CDW}$ and $S_{\rm SC}$, respectively. In stark contrast, both the CDW and SC structure factors exhibit significant finite-size dependence, suggesting their coexistence in this regime.
Figure 9: (a) Charge and pair structure factors as a function of the phonon frequency $\omega_{0}$ at fixed $\beta=L=10$ and $\lambda/\alpha=6$. The horizontal dashed line marks the corresponding degenerate $S_{\rm CDW}$ and $S_{\rm SC}$ value in the antiadiabatic limit $(\omega_{0}\to\infty)$ for the same parameters. (b) Pair structure factor as a function of $\omega_{0}$ for several lattice sizes $L$, with fixed $\beta=L$ and $\lambda/\alpha=6$. Inset: extrapolation of $S_{\rm sc}/L^{2}$ to the thermodynamic limit, for fixed $\omega_{0}=48$. For both panels, we set $t=0$ and $\alpha=1$.
The previous analysis suggests that any CDW-SC coexistence at finite $\omega_{0}$ would be restricted to large values of $\omega_{0}$, in the regime of pure Rashba hopping. To estimate the required scale, Fig. 9 (a) compares $S_{\mathrm{CDW}}$ and $S_{\mathrm{SC}}$ as functions of $\omega_{0}$. Both quantities trend toward the attractive Hubbard limit (horizontal dashed line) and become degenerate with increasing frequency. However, notice that achieving (results quantitatively consistent with) the antiadiabatic limit requires $\omega_{0}/\alpha\gtrsim 64$, i.e. $\omega_{0}$ roughly an order of magnitude larger than the bandwidth. With this in mind, we turn to examine the behavior of the SC order parameter at large frequencies. Figure 9 (b) shows $S_{\mathrm{SC}}$ as a function of $\omega_{0}$ for several lattice sizes. Notice that, for $\omega_{0}/\alpha\gtrsim 24$, $S_{\mathrm{SC}}$ increases with system size, which is consistent with the occurrence of long-range order. Extrapolating these data for large frequencies to the thermodynamic limit leads a nonzero SC order parameter (as showed in inset). This confirms coexistence of SC and CDW order at finite $\omega_{0}$, while the two become degenerate (imposed by symmetry) in the limit of $\omega_{0}\to\infty$ [^1] We expect that a similar behavior should occur even for $t/\alpha\neq 0$, since both non-interacting particle-hole and particle-particle channels diverges logarithmically..
@@ IV Conclusions
However, the features of the model change drastically in the limit of a pure Rashba hopping, i.e., when $t/\alpha=0$. In this limit, four Weyl cones appear at the half-filling, and quantum phase transitions are expected to occur at strong interactions. Specifically, we demonstrated that the CDW phase emerges for a finite quantum critical coupling $\lambda_{c}$, for a given phonon frequency. By varying $\omega_{0}$, the critical $\lambda_{c}$ also changes, leading to the phase diagram shown in Fig. 6. At finite phonon frequencies, we expect this transition to belong to the Ising Gross-Neveu universality class with $N_{f}=8$ two component Weyl cones Herbut et al. (2009); Huffman and Chandrasekharan (2020). In the antiadiabatic limit where $\omega_{0}\to\infty$ the interaction corresponds to the attractive Hubbard model, with $|U|=\lambda$. In this limit, and as described in Sec. II.2 we expect an emergent SU(2)η symmetry the ties SC and CDW orders together. For large frequencies ($\omega_{0}\gtrsim 24$), SC emerges, with the ground state exhibiting a coexistence between CDW and SC. These quantities become degenerate when $\omega_{0}\to\infty$. This corresponds to the same universality class as that of the Hubbard model on the honeycomb lattice Assaad and Herbut (2013); Otsuka et al. (2016). These findings provide a step forward in understanding the interplay between SC and CDW phases in systems with spin-orbit coupling, and we hope they could shed light on the nature of these compounds.
@@ ACKNOWLEDGMENTS
The authors are grateful to the Brazilian Agencies Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Ensino Superior (CAPES), and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro, FAPERJ. J.F. thanks FAPERJ, Grant No. SEI-260003/019642/2022, and ANID Fondecyt grant No. 3240320. S.A.S.J. thanks CNPq, grant No. 201000/2024-5. N.C.C. acknowledges support from FAPERJ Grants No. E-26/200.258/2023 [SEI-260003/000623/2023] and E-26/210.592/2025 [SEI-260003/004500/2025], CNPq Grants No. 313065/2021-7 and 308130/2025-1, and Serrapilheira Institute Grant No. R-2502-52037. FA thanks the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490). This research was partially supported by the supercomputing infrastructure of the NLHPC (Grant No. CCSS210001).